OpenStudy (anonymous):

Which of the points is not a solution to x-2y=4 and my answers are(0,-2) , (0,2) , (2,-1), (4,0)

OpenStudy (anonymous):

B and D are not solutions..

OpenStudy (anonymous):

Plug each point into the equation and see if they're true. For example, take (0, -2) and plug it in like this: 0-2(-2)=4 --> 4=4, so (0, -2) is a solution.

OpenStudy (anonymous):

Still didn't help...

OpenStudy (anonymous):

sorry, i meant B and C are not solutions..:)

OpenStudy (anonymous):

I just need one

OpenStudy (anonymous):

C is a solution. 2-2(-1)=4. 2+2=4. The only one that isn't a solution is B.

OpenStudy (anonymous):

Plz help me

OpenStudy (anonymous):

ooh.. my bad.. @girlnotonfire is right..:)

OpenStudy (anonymous):

@Hailey553 I thought you were right before! It seemed so simple and yet it confused me :P

OpenStudy (anonymous):

:)